# LENS: Learning to Segment Anything with Unified Reinforced Reasoning

### v1

@@ Abstract
Text-prompted image segmentation enables fine-grained visual understanding and is critical for applications such as human-computer interaction and robotics. However, existing supervised fine-tuning methods typically ignore explicit chain-of-thought (CoT) reasoning at test time, which limits their ability to generalize to unseen prompts and domains. To address this issue, we introduce LENS, a scalable reinforcement-learning framework that jointly optimizes the reasoning process and segmentation in an end-to-end manner. We propose unified reinforcement-learning rewards that span sentence-, box-, and segment-level cues, encouraging the model to generate informative CoT rationales while refining mask quality. Using a publicly available 3-billion-parameter vision–language model, i.e., Qwen2.5-VL-3B-Instruct, LENS achieves an average cIoU of 81.2% on the RefCOCO, RefCOCO+, and RefCOCOg benchmarks, outperforming the strong fine-tuned method, i.e., GLaMM, by up to 5.6%. These results demonstrate that RL-driven CoT reasoning serves as a robust prior for text-prompted segmentation and offers a practical path toward more generalizable Segment Anything models.
@@ Introduction
Recent studies (Lai et al. 2024; Bai et al. 2024; Ren et al. 2024) incorporate multimodal large language models (MLLMs) to improve cross-modal localization. These methods drive the segmentation process with a single token “<seg>” and train the entire pipeline via supervised fine-tuning (SFT). However, this paradigm faces two major limitations: (i) it neglects the intermediate reasoning process that is essential for complex, reasoning-intensive tasks, and (ii) its heavy reliance on SFT often leads to overfitting and weak generalization. These issues motivate us to pursue a more robust and generalizable test-time reasoning framework.
@@
Group Relative Policy Optimization (GRPO) (Shao et al. 2024) is a rule-based reinforcement-learning (RL) algorithm for post-training large language models. By optimizing policies with group-relative rewards, GRPO strengthens reasoning ability and empirically generalizes better than SFT. Motivated by GRPO, we present LENS, a unified, test-time reasoning framework for text-prompted segmentation. LENS adopts the GRPO strategy with the proposed unified rewards that simultaneously consider sentence-, box-, and segment-level cues. Furthermore, we introduce a context module to bridge the MLLM and the segmentation model, which extracts the reasoning and grounding information as prior to guide the segmentation process.
Concurrent work SegZero (Liu et al. 2025) directly feeds the grounding output, i.e., the bounding box and points, of an MLLM into a frozen SAM, achieving strong reasoning segmentation ability. However, the final mask quality is fully determined by the MLLM’s grounding, and any errors in grounding will propagate downstream, leading to segmentation errors. In contrast, LENS offers an end-to-end solution that jointly optimizes language understanding and pixel-wise mask prediction (Fig. 1). We establish a tight coupling between the MLLM and SAM through the proposed context module and a pre-alignment stage. Moreover, the proposed unified reinforcement learning rewards, i.e., format reward, box-IoU reward, and segment-IoU reward, encourage the model to produce informative chain-of-thought rationales while refining mask quality. Extensive experiments show that LENS sets a new state-of-the-art on standard text-prompted segmentation benchmarks.
@@
  • LENS achieves 81.2% average cIoU on RefCOCO, RefCOCO+, and RefCOCOg, 58.0% cIoU on ReasonSeg-Test, and 78.3% cIoU on GS-Eval, establishing new state-of-the-art performance for text-prompted segmentation.
@@
Text-prompted segmentation, also known as referring segmentation, aims to segment objects described by natural language expressions. This task requires both visual understanding and language comprehension capabilities. Early work includes ReferIt (Kazemzadeh et al. 2014) dataset and corresponding models that localize objects based on natural language descriptions.
Recent advances have focused on multimodal fusion architectures. LAVT (Yang et al. 2022) proposed a lightweight vision-transformer approach for referring segmentation. RefTR (Li and Sigal 2021) employed transformer-based cross-modal fusion. CRIS (Wang et al. 2022) introduced contrastive learning for improved text-image alignment. More recent work includes X-Decoder (Zou et al. 2023a) and SEEM (Zou et al. 2023b), which unify various segmentation tasks including text-prompted segmentation.
The integration of large language models has opened new possibilities. LISA (Lai et al. 2024) combines SAM with large language models for reasoning segmentation. PerSAM (Zhang et al. 2023) enables few-shot personalization of SAM. However, these approaches primarily rely on supervised fine-tuning, which may limit their reasoning capabilities and generalization to unseen scenarios.
@@
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone technique for aligning large language models with human preferences. PPO (Proximal Policy Optimization) (Schulman et al. 2017a) serves as the foundational algorithm, enabling stable policy updates while preventing catastrophic forgetting.
Recent breakthroughs include InstructGPT (Ouyang et al. 2022) and ChatGPT, which demonstrated the effectiveness of RLHF in producing helpful and harmless responses. Constitutional AI (Bai et al. 2022) proposed training models to follow a set of principles. DPO (Direct Preference Optimization) (Rafailov et al. 2023) simplified the RLHF pipeline by directly optimizing on preference data.
The emergence of reasoning-focused RL approaches has shown remarkable success. DeepSeek-R1 (Guo et al. 2025) demonstrated that large-scale reinforcement learning can significantly improve reasoning capabilities, achieving competitive performance with leading models. Chain-of-thought reasoning with RL supervision has been explored in various contexts (Zelikman et al. 2022; Huang et al. 2022), showing that explicit reasoning processes can be learned and improved through reinforcement learning.
@@
In this section, we introduce the proposed LENS architecture and the proposed reinforcement learning training strategy, as shown in Fig. 2. First, we introduce the proposed LENS architecture that enables end-to-end reasoning and segmentation. Then, we introduce the proposed pretraining alignment stage that establishes the foundational connection between the MLLM and SAM. Finally, we introduce the proposed reinforcement learning stage that jointly optimizes the model’s reasoning and segmentation capabilities.
@@
The concurrent SegZero (Liu et al. 2025) triggers a frozen SAM with bounding boxes and points predicted by an MLLM, which achieves strong reasoning ability. However, its non-end-to-end nature hinders the joint alignment of understanding and segmentation representations and ultimately yields sub-optimal segmentation masks. To overcome this limitation, we introduce LENS, an end-to-end reinforcement-learning framework that couples an MLLM with SAM through a context module. Thereby, the model can jointly optimize the reasoning and segmentation capabilities under the multi-grained unified rewards and pixel-level entropy loss.
@@
We use Multi-Modal Large Language Models (MLLMs), i.e., Qwen2.5-VL (Bai et al. 2025), as the reasoning model $\pi_{\text{mllm}}$. Previous MLLMs (Liu et al. 2023; Bai et al. 2023) show promising performance in object-centric localization, but still lack pixel-level perception ability. Some work (Lai et al. 2024; Bai et al. 2024) incorporates pre-trained segmentation models into MLLMs to compensate the lack of pixel-level perception ability. However, these methods simply use a segmentation token in short response, i.e., ‘It is $\mathtt{<seg>}$’, to guide the segmentation, which have neglected the reasoning process, e.g., Chain-of-Thought (CoT) (Wei et al. 2022). To address this, we propose to utilize the reasoning process to guide the segmentation, which brings more robust and effective segmentation ability. We input the system prompt $p_{\text{sys}}$, textual instruction $T$ and image $I$ to the reasoning model, and obtain the CoT $P_{\text{CoT}}$ and box prediction $P_{\text{box}}$.
@@
Considering concurrent method (Liu et al. 2025) only rely detokenized bounding box and points to prompt the SAM, which leads to suboptimal performance and eliminates the end-to-end optimization. To address this, we propose a context module to bridge the gap between understanding and segmentation during the reinforcement learning process. The context module contains the context query and the connector. After the MLLM generates the CoT and box prediction, we use the context query to extract the context information from the MLLM. Then, we use the connector to project the context information as the segmentation prompt for the SAM.
@@
Nowadays, SAM (Kirillov et al. 2023) is the most popular segmentation model, which accepts prompts including point and box and produces precise segmentation results. Some methods (Zhang et al. 2024b; Lai et al. 2024) incorporate MLLMs, which enable SAM to segment objects with text prompts. However, these methods output the single segmentation token $\mathtt{<seg>}$’ directly, which neglects the reasoning process and token capacity. To address this, we propose to use multiple queries to extract reasoning context and ensure the enough token capacity. Moreover, we incorporate pixel-level entropy loss to reinforcement learning training, which provides more fine-grained optimization guidance. We feed the projected segmentation prompt $Q_{\text{seg}}$ to the SAM model $\pi_{\text{sam}}$ for segmentation $P_{\text{seg}}$.
@@
The loss function during this stage is purely based on segmentation objectives:
@@
While the pretraining alignment stage enables basic multimodal understanding and segmentation, the quality of reasoning process significantly impacts segmentation performance. Existing methods (Liu et al. 2025; Shen et al. 2025) typically optimize only the understanding part, overlooking the crucial segmentation decoder.
@@
The group relative policy optimization (GRPO) (Shao et al. 2024) introduces a new rule-based reinforcement learning algorithm for post-training optimization of large language models. Traditional reinforcement learning algorithms, i.e., PPO (Schulman et al. 2017b), use separate critic models to evaluate the quality of the policy model, while leading to high memory and computational costs. To alleviate this problem, GRPO eliminates the separate critic model and only uses group relative rewards to guide the optimization of the policy model. Given a question sample $q$, GRPO first samples a group of outputs $\{o_{i}\}_{i=1}^{G}$ with the policy model $\pi$, where $G$ is the number of samples. Next, it uses reward function $\mathtt{R}(\cdot)$ to calculate the rewards for each output. To determine the relative quality of the outputs in the group, GRPO then calculates the advantage $A_{i}$ for each output as follows:

### v2

@@ Abstract
Text-prompted image segmentation enables fine-grained visual understanding and is critical for applications such as human-computer interaction and robotics. However, existing supervised fine-tuning methods typically ignore explicit chain-of-thought (CoT) reasoning at test time, which limits their ability to generalize to unseen prompts and domains. To address this issue, we introduce LENS, a scalable reinforcement-learning framework that jointly optimizes the reasoning process and segmentation in an end-to-end manner. We propose unified reinforcement-learning rewards that span sentence-, box-, and segment-level cues, encouraging the model to generate informative CoT rationales while refining mask quality. Using a publicly available 3-billion-parameter vision–language model, i.e., Qwen2.5-VL-3B-Instruct, LENS achieves an average cIoU of 81.2% on the RefCOCO, RefCOCO+, and RefCOCOg benchmarks, outperforming the strong fine-tuned method, i.e., GLaMM, by up to 5.6%. These results demonstrate that RL-driven CoT reasoning significantly enhances text-prompted segmentation and offers a practical path toward more generalizable Segment Anything models (SAM).
@@ Introduction
Recent studies (lai2024lisa; bai2024onetokenseg; ren2024pixellm) incorporate multimodal large language models (MLLMs) to improve cross-modal localization. These methods drive the segmentation process with a single token “<seg>” and train the entire pipeline via supervised fine-tuning (SFT). However, this paradigm faces two major limitations: (i) it neglects the intermediate reasoning process that is essential for complex, reasoning-intensive tasks, and (ii) its heavy reliance on SFT often leads to overfitting and weak generalization. These issues motivate us to pursue a more robust and generalizable test-time reasoning framework.
@@
Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) is a rule-based reinforcement-learning (RL) algorithm for post-training large language models. By optimizing policies with group-relative rewards, GRPO strengthens reasoning ability and empirically generalizes better than SFT. Motivated by GRPO, we present LENS, a unified, test-time reasoning framework for text-prompted segmentation. LENS adopts the GRPO strategy with the proposed unified rewards that simultaneously consider sentence-, box-, and segment-level cues. Furthermore, we introduce a context module to bridge the MLLM and the segmentation model, which extracts the reasoning and grounding information as prior to guide the segmentation process.
Concurrent work SegZero (liu2025segzero) directly feeds the grounding output, i.e., the bounding box and points, of an MLLM into a frozen SAM, achieving strong reasoning segmentation ability. However, the final mask quality is fully determined by the MLLM’s grounding, and any errors in grounding will propagate downstream, leading to segmentation errors. In contrast, LENS offers an end-to-end solution that jointly optimizes language understanding and pixel-wise mask prediction (Fig. 1). We establish a tight coupling between the MLLM and SAM through the proposed context module and a pre-alignment stage. Moreover, the proposed unified reinforcement learning rewards, i.e., format reward, box-IoU reward, and segment-IoU reward, encourage the model to produce informative chain-of-thought rationales while refining mask quality. Extensive experiments show that LENS sets a new state-of-the-art on standard text-prompted segmentation benchmarks.
@@
  • LENS achieves 81.2% average cIoU on RefCOCO-series benchmarks, 58.0% cIoU on ReasonSeg-Test, and 78.3% cIoU on GS-Eval, establishing new state-of-the-art performance for text-prompted segmentation.
@@
Text-prompted segmentation, i.e., referring segmentation, aims to segment objects described by natural language expressions. This task requires both visual understanding and language comprehension capabilities. Early work includes ReferIt (kazemzadeh2014referitgame) dataset and corresponding models that localize objects based on natural language descriptions.
Recent advances have focused on multimodal fusion architectures. LAVT (yang2022lavt) proposed a lightweight vision-transformer approach for referring segmentation. RefTR (li2021referring) employed transformer-based cross-modal fusion. CRIS (wang2022cris) introduced contrastive learning for improved text-image alignment. More recent work includes X-Decoder (zou2023generalized) and SEEM (zou2023segment), which unify various segmentation tasks including text-prompted segmentation.
The integration of large language models has opened new possibilities. LISA (lai2024lisa) combines SAM with large language models for reasoning segmentation. PerSAM (zhang2023personalize) enables few-shot personalization of SAM. However, these approaches primarily rely on supervised fine-tuning, which may limit their reasoning capabilities and generalization to unseen scenarios.
@@
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone technique for aligning large language models with human preferences. PPO (Proximal Policy Optimization) (schulman2017proximal) serves as the foundational algorithm, enabling stable policy updates while preventing catastrophic forgetting.
Recent breakthroughs include InstructGPT (ouyang2022training) and ChatGPT, which demonstrated the effectiveness of RLHF in producing helpful and harmless responses. Constitutional AI (bai2022constitutional) proposed training models to follow a set of principles. DPO (Direct Preference Optimization) (rafailov2023direct) simplified the RLHF pipeline by directly optimizing on preference data.
The emergence of reasoning-focused RL approaches has shown remarkable success. DeepSeek-R1 (guo2025deepseek) demonstrated that large-scale reinforcement learning can significantly improve reasoning capabilities, achieving competitive performance with leading models. Chain-of-thought reasoning with RL supervision has been explored in various contexts (zelikman2022star; huang2022large), showing that explicit reasoning processes can be learned and improved through reinforcement learning.
@@
In this section, we introduce the proposed LENS architecture and the proposed reinforcement learning training strategy. First, we introduce the proposed LENS architecture that enables end-to-end reasoning and segmentation. Then, we introduce the proposed pretraining alignment stage that establishes the foundational connection between the MLLM and SAM. Finally, we introduce the proposed reinforcement learning stage that jointly optimizes the model’s reasoning and segmentation capabilities.
@@
The concurrent SegZero (liu2025segzero) triggers a frozen SAM with bounding boxes and points predicted by an MLLM, which achieves strong reasoning ability. However, its non-end-to-end nature hinders the joint alignment of understanding and segmentation representations and ultimately yields sub-optimal segmentation masks. To overcome this limitation, we introduce LENS, an end-to-end reinforcement-learning framework that couples an MLLM with SAM through a context module. Thereby, the model can jointly optimize the reasoning and segmentation capabilities under the multi-grained unified rewards and pixel-level entropy loss.
@@
We use MLLMs, i.e., Qwen2.5-VL (bai2025qwen2p5vl), as the reasoning model $\pi_{\text{mllm}}$. Previous MLLMs (liu2023llava; bai2023qwenvlversatilevisionlanguagemodel) show promising performance in object-centric localization, but still lack pixel-level perception ability. Some work (lai2024lisa; bai2024onetokenseg) incorporates pre-trained segmentation models into MLLMs to compensate the lack of pixel-level perception ability. However, these methods simply use a segmentation token in short response, i.e., ‘It is $\mathtt{<seg>}$’, to guide the segmentation, which have neglected the reasoning process, e.g., Chain-of-Thought (CoT) (wei2022cot). To address this, we propose to utilize the reasoning process to guide the segmentation, which brings more robust and effective segmentation ability. We input the system prompt $p_{\text{sys}}$, textual instruction $T$ and image $I$ to the reasoning model, and obtain the CoT $P_{\text{CoT}}$ and box prediction $P_{\text{box}}$.
@@
Concurrent method (liu2025segzero) only rely detokenized bounding box and points to prompt the SAM, which leads to suboptimal performance and eliminates the end-to-end optimization. To address this, we propose a context module to bridge the gap between understanding and segmentation during the reinforcement learning process. The context module contains the context query and the connector. After the MLLM generates the CoT and box prediction, we use the context query to extract the context information from the MLLM. Then, we use the connector to project the context information as the prompt.
@@
SAM (kirillov2023sam) is the most popular segmentation model (cheng2025cross), which accepts prompts including point and box and produces precise segmentation results. Some methods (zhang2024evfsam; lai2024lisa) incorporate MLLMs, which enable SAM to segment objects with text prompts. However, these methods output the single segmentation token $\mathtt{<seg>}$’ directly, which neglects the reasoning process and token capacity. To address this, we propose to use multiple queries to extract reasoning context and ensure the enough token capacity. Moreover, we incorporate pixel-level entropy loss to reinforcement learning training, which provides more fine-grained optimization guidance. We feed the projected segmentation prompt $Q_{\text{seg}}$ to the SAM model $\pi_{\text{sam}}$ for segmentation $P_{\text{seg}}$.
@@
We use segmentation objectives in this stage:
@@
While the pretraining alignment stage enables basic multimodal understanding and segmentation, the quality of reasoning process significantly impacts segmentation performance. Existing methods (liu2025segzero; shen2025vlmr1) typically optimize only the understanding part, overlooking the crucial segmentation decoder.
@@
The GRPO (shao2024deepseekmath) introduces a new rule-based reinforcement learning algorithm for post-training optimization of large language models. Traditional reinforcement learning algorithms, i.e., PPO (schulman2017ppo), use separate critic models to evaluate the quality of the policy model, while leading to high memory and computational costs. To alleviate this problem, GRPO eliminates the separate critic model and only uses group relative rewards to guide the optimization of the policy model. GRPO employs reward function $R(\cdot)$ to evaluate the outputs and Kullback-Leibler (KL) divergence penalty $\mathbb{D}_{\text{KL}}$ to regularize the policy update. Due to the page limitation, we leave more details in the appendix.
Method RefCOCO RefCOCO+ RefCOCOg AVG
val testA testB val testA testB val test
without active chain-of-thought (CoT) reasoning
LAVT (yang2022lavt)(CVPR 22) 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1 65.8
ReLA (liu2023gres)(CVPR 23) 73.8 76.5 70.2 66.0 71.0 57.7 65.0 66.0 68.3
X-Decoder (zou2023generalized)(CVPR 23) - - - - - - 64.6 - -
SEEM (zou2023segment)(NeurIPS 23) - - - - - - 65.7 - -
LISA (lai2024lisa) (CVPR 24) 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5 67.9
PixelLM (ren2024pixellm)(CVPR 24) 76.9 78.5 74.4 69.2 72.1 64.5 70.7 72.4 72.3
PerceptionGPT-7B (pi2024perceptiongpt)(CVPR 24) 75.1 78.6 71.7 68.5 73.9 61.3 70.3 71.7 71.4
OMG-LLaVA (zhang2024omg)(NeurIPS 24) 78.0 80.3 74.1 69.1 73.1 63.0 72.9 72.9 72.9
VISA (yan2024visa)(ECCV 24) 72.4 75.5 68.1 59.8 64.8 53.1 65.5 66.4 65.7
GLaMM (rasheed2024glamm)(CVPR 24) 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9 75.6
SAM3-Agent-Gemini2.5-Pro (OpenReview 2025) 74.9 77.8 69.9 66.9 71.1 62.4 73.3 73.6 74.2
with active chain-of-thought (CoT) reasoning
Seg-Zero-3B (liu2025segzero) (arXiv 25) - 79.3 - - 73.7 - 71.5 - -
Seg-Zero-7B (liu2025segzero) (arXiv 25) - 80.3 - - 76.2 - 72.6 - -
LENS-2B (Ours) 80.7 82.7 77.1 73.8 78.0 67.3 75.7 76.4 76.5
LENS-3B (Ours) 84.2 85.3 81.0 79.4 82.8 74.3 81.2 81.0 81.2
Table 1: Comparison on the Referring Expression Segmentation (RES) task using cIoU as the evaluation metric. Our method achieves state-of-the-art performance across all RefCOCO, RefCOCO+, and RefCOCOg benchmarks.Bold indicates the best (SoTA) performance, and underlined denotes the second-best. The “active chain-of-thought (CoT) reasoning” means using explicit chain-of-thought at test time.
GRPO with Unified Rewards.
Unlike standard VLM-R1 methods (shen2025vlmr1; yu2025perceptionr1) that relies on purely understanding reward signal, i.e., format reward and box IoU reward, we introduce a comprehensive multi-faceted reward system specifically designed for both understanding and segmentation. Our reward function encompasses three complementary components:
  • Box IoU Reward ($R_{\text{box}}$): Measures localization accuracy through IoU between predicted and ground truth bounding boxes. The range is $[0,1]$.
The unified reward function is formulated as:
where $\{\lambda_{i}\}_{i=1}^{3}$ are balancing hyperparameters.
Training Objective.
$\mathcal{J}_{\text{LENS}}(\theta)$ $=\mathcal{J}(\theta;R_{\text{unified}},\mathbb{D}_{\mathrm{KL}})+\alpha\mathcal{L}_{\text{seg}}(P_{\text{seg}},M_{\text{gt}})$ (7)
Experiment
Implementation Details
We employ Qwen-VL series models(bai2025qwen2p5vl; wang2024qwen2) as the reasoning model and SAM2-Large (ravi2024sam2segmentimages) as the segmentation model. Specifically, LENS-2B is utilizes Qwen2-VL and LENS-3B uses Qwen2.5-VL. Training is conducted on 16 NVIDIA L40S GPUs, with our pipeline built upon the DeepSpeed engine (rasley2020deepspeed). Detailed experimental settings are summarized in the Appendix.
Benchmarks.
We evaluate the proposed framework on multiple benchmarks: Referring Expression Segmentation (RES) benchmarks (RefCOCO series) (nagaraja2016modeling; yu2016modeling), GroundingSuite benchmarks (hu2025groundingsuite), and Reasoning Segmentation benchmarks (lai2024lisa). The quantitative comparisons are presented in Table 1 and Table 2, respectively.
Evaluation Metrics.
Following previous works on referring segmentation (kazemzadeh2014referitgame), we adopt two commonly used evaluation metrics: generalized IoU (gIoU) and cumulative IoU (cIoU). Specifically, gIoU is computed as the average of per-image Intersection-over-Union (IoU) scores, while cIoU is defined as the IoU of the cumulative predicted and ground-truth masks across the entire dataset.
Main Results
Method ReasonSeg-Val ReasonSeg-Test GroundingSuite-Eval
gIoU cIoU gIoU cIoU gIoU cIoU
PSALM (zhang2024psalm)(ECCV 24) - - - - 39.2 34.4
[…]
LISA-7B(ft) (lai2024lisa) (CVPR 24) 52.9 54.0 55.6 56.9 60.9 68.6
Seg-Zero-3B (liu2025segzero) (arXiv 25) 58.2 53.1 56.1 48.6 67.3 68.9
SAM3-Agent-Qwen2.5-VL-7B (OpenReview 2025) 65.4 50.5 62.6 56.2 - -
LENS-3B (Ours) 62.1 64.9 57.2 58.0 67.0 78.3
Table 2: Performance comparison on the ReasonSeg and GroundingSuite-Eval benchmarks, evaluated using gIoU and cIoU. Bold and underlined entries indicate the best and second-best results, respectively. $\ddagger$ marks models trained on both train and test splits of ReasonSeg. All models are not trained on GroundingSuite. LENS shows superior reasoning and generalization ability.
Note that PSALM (zhang2024psalm), HyperSeg (wei2024hyperseg), and InstructSeg (wei2024instructseg) utilize pre-trained Mask2Former (cheng2022mask2former) models, whose training data, i.e., COCO 2017 (lin2014microsoft), has overlap with the RefCOCO benchmarks. Therefore, we compare with them in Table 2, evaluating their performance on the Reasoning Segmentation and GroundingSuite-Eval benchmarks.
Query Num. 1 16 32 64 128
cIoU (%) 85.9 87.4 87.8 87.9 87.2
Connector MLP ViT
cIoU (%) 72.8 87.8
Context Module Align. Stage RL Stage cIoU
$R_{\text{box}}$ $R_{\text{format}}$ $R_{\text{seg}}$
$\times$ $\times$ $\times$ $\times$ $\times$ 79.3
$\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
$\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ 80.9
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ 72.3
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 81.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 82.7
Table 5: Ablation study of components on RefCOCO testA.
Referring Expression Segmentation.
As shown in Table 1, the proposed LENS achieves state-of-the-art performance across all splits of the RefCOCO, RefCOCO+, and RefCOCOg benchmarks. Compared to prior models such as GlaMM (rasheed2024glamm), our method delivers significant improvements, especially on RefCOCO+ and RefCOCOg, which are known to require a more precise understanding of compositional and spatial language. Notably, LENS-3B surpasses the best previous method by 6.1% on RefCOCOg-test, highlighting the model’s strength in handling complex referring expressions.
ReasonSeg and GroundingSuite-Eval Benchmarks.
As shown in Table 2, we evaluate the proposed LENS on the ReasonSeg and GroundingSuite-Eval benchmarks. GroundingSuite (GS) is designed for challenging scenarios, and none of the methods in Table 2 are trained on GS-Train. On the zero-shot GS-Eval benchmark, LENS achieves a 9.4% improvement over the second-best method, demonstrating strong out-of-domain (OOD) generalization.
In terms of segmentation quality, LENS achieves competitive results in gIoU and exhibits a clear lead in cIoU, which captures holistic segmentation performance across objects and scenes. For example, on GS-Eval, LENS attains a cIoU of 78.3%, surpassing Seg-Zero-3B by 9.4%.
Ablation Studies
Number of Queries.
Connector Design.
We further investigate the role of the proposed connector bridging the MLLM and the segmentation head. The connector is implemented as a 4-layer Qwen-2.5 Transformer with a hidden dimension of 2048, matching the MLLM output. As shown in Table 4, replacing a simple MLP with this Transformer significantly improves performance. Overall, these studies underscore the importance of architectural choices in enabling effective reasoning and accurate segmentation within the proposed unified framework.
Framework Analysis.
To investigate the contribution of each component in our framework, we conduct an ablation study on the RefCOCO testA split, as shown in Table 5. We start with a Lisa-liked baseline (lai2024lisa) model without the context module, alignment stage, or reinforcement learning (RL), which achieves a cIoU of 79.3%. Directly adding the context module and applying the box reward lead to instable. After enabling format reward, the model captures basic localization cues, but the lack of explicit alignment between textual and visual features leads to sub-optimal performance. To address this issue, we enable the alignment stage, where the context module is pretrained to align MLLM and SAM. This significantly improves the segmentation accuracy, boosting cIoU to 81.9%, and demonstrates the critical importance of semantic alignment in the proposed unified framework. Finally, introducing the segmentation-level reward ($R_{\text{seg}}$) further refines the mask quality, pushing the performance to 82.7% cIoU.
Visualizations
We provide qualitative results to illustrate the effectiveness of our framework on both general and reasoning-intensive segmentation tasks. Fig. 3 shows examples from the RefCOCOg benchmark. LENS accurately segments target objects based on referring expressions, demonstrating strong performance in scenes with partially obscure objects and multiple similar instances. Notably, the proposed unified framework and end-to-end optimization make LENS robust to correct the potential error in the box prior. Fig. 4 presents results on the reasoning segmentation benchmark. LENS exhibits robust test-time reasoning capabilities, effectively handling tasks that require spatial understanding, multi-step inference, and comparative reasoning. These visualizations highlight the LENS’s ability to integrate high-level textual reasoning with fine-grained visual perception.
Conclusion
Acknowledgement: This work was partially supported by the National Natural Science Foundation of China (No. 62276108).
Appendix A Appendices
More GRPO Details
Given a question sample $q$, GRPO first samples a group of outputs $\{o_{i}\}_{i=1}^{G}$ with the policy model $\pi$, where $G$ is the number of samples. Next, it uses reward function $\mathtt{R}(\cdot)$ to calculate the rewards for each output. To determine the relative quality of the outputs in the group, GRPO then calculates the advantage $A_{i}$ for each output as follows: